2. Three cards are drawn from a standard deck w/out replacing them.
a.) what is the probability that they are all face cards?
b.) what is the probability that they are all from the same suit?
3. A pair of fair dice are rolled. Write down 6 by 6 sample space and find the probability that the sum of the #'s showing is 8 or 9.
4. 40% of the students in a school take English, 30% take math, and 25% take both math & English. A student is picked at random, what is the probability that they take Math or English?

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Three cards are drawn from a standard deck w/out replacing them.
a.) what is the probability that they are all face cards?
Ans: 12C3/52C3 = 0.0100
------------------------------------------
b.) what is the probability that they are all from the same suit?
Ans: 4*13C3/52C3
-----------------------------
3. A pair of fair dice are rolled. Write down 6 by 6 sample space and find the probability that the sum of the #'s showing is 8 or 9.
8's:: 5 combinations
9's:: 4 combinations
Ans: (5+4)/36 = 1/4
--------------------------------
4. 40% of the students in a school take English, 30% take math, and 25% take both math & English. A student is picked at random, what is the probability that they take Math or English?
P(m OR e) = P(m)+P(e)-P(m AND e) = 0.3+0.4-0.25 = 0.45
-----------------------------------
Cheers,
Stan H.
--------------

Answer by addingup(3677)

4suits, each has 3 face cards: 4 x 3 = 12
Since we are not replacing:
first draw: 12 possible events out of a total of 52
second : 11 out of 51
third : 10 out of 50
P(f*f*f) = 12/52*11/51*10/50 = 11/1105 = 0.01 or 1%
:
John